OpenStudy (boxman61):

solve using quadratic formula.. 5u^2-2u+1=0

OpenStudy (anonymous):

Do you know what the quadratic formula is?

OpenStudy (boxman61):

i have it broken down to \[2\pm \sqrt{-16}/10\]

OpenStudy (anonymous):

That looks correct

OpenStudy (boxman61):

yes i do.... im just confused with the simplifying part of this equation...answer is \[1/5\pm2/5i\]

OpenStudy (boxman61):

sure, i actually got it by \[2\pm\sqrt{4-20}/10\]

OpenStudy (anonymous):

\[\sqrt{-16} = \sqrt{-1 * 16} = \sqrt{-1} * \sqrt{16}\] That's true right?

OpenStudy (boxman61):

yes

OpenStudy (anonymous):

So you simplify the root 16 to 4, and sqrt -1 to i

OpenStudy (anonymous):

And you get something like \[(2 +/- 4i ) / 10\]

OpenStudy (anonymous):

Then you just take it one step more, and simplify the divide by 10 to each term?

OpenStudy (boxman61):

thanks

OpenStudy (anonymous):

No problem